# Train Once, Deploy Anywhere: Realize Data-Efficient Dynamic Object Manipulation

### v1

@@ Abstract
Realizing generalizable dynamic object manipulation is important for enhancing manufacturing efficiency, as it eliminates specialized engineering for various scenarios. To this end, imitation learning emerges as a promising paradigm, leveraging expert demonstrations to teach a policy manipulation skills. Although the generalization of an imitation learning policy can be improved by increasing demonstrations, demonstration collection is labor-intensive. To address this problem, this paper investigates whether strong generalization in dynamic object manipulation is achievable with only a few demonstrations. Specifically, we develop an entropy-based theoretical framework to quantify the optimization of imitation learning. Based on this framework, we propose a system named Generalizable Entropy-based Manipulation (GEM). Extensive experiments in simulated and real tasks demonstrate that GEM can generalize across diverse environment backgrounds, robot embodiments, motion dynamics, and object geometries. Notably, GEM has been deployed in a real canteen for tableware collection. Without any in-scene demonstration, it achieves a success rate of over 97% across more than 10,000 operations.
Index Terms:
Dynamic object manipulation, imitation learning, information entropy, manufacturing efficiency.
@@ I Introduction
Robotic manipulation of dynamic objects has broad applications in industrial manufacturing, particularly in conveyor-based assembly lines where robots handle transported products [1, 2]. Compared with the extensively studied static object manipulation problem [3, 4], dynamic object manipulation poses more challenges due to its demand for more time-efficient motion planning and stricter operational precision. Addressing these challenges is important for enhancing the efficiency and flexibility of industrial automation systems, thereby enabling more economical manufacturing.
@@
Therefore, although realizing generalizable dynamic object manipulation is of huge importance, collecting demonstrations is costly and there are few public data resources [13]. Given these facts, this work studies whether strong generalization in dynamic object manipulation can be achieved without relying on numerous demonstrations. To address this question, we delve into the optimization essence of imitation learning, framing it as a distribution approximation problem. Building upon this formulation, we propose an information entropy based theoretical framework to quantify the amount of data required to achieve generalizable manipulation. From this theoretical framework, we identify two key aspects: first, reducing the complexity of redundant information in observations, and second, decreasing the ambiguity in action prediction.
@@
Figure 1: Primarily using demonstrations collected from a simulator, our method can generalize across diverse environment backgrounds, robot embodiments, motion dynamics, and object geometries. Our method has been reliably deployed in a real canteen to conduct tableware collection. Without relying on any demonstration collected in this canteen, our method achieves a success rate of over 97% in seven consecutive days of operation, performing more than 10,000 tableware collection operations.
According to these two key aspects, we develop a manipulation system named Generalizable Entropy-based Manipulation (GEM). GEM consists of two blocks, i.e., object-centric geometry perception and hybrid action control. In object-centric geometry perception, we devise a representation that emphasizes the geometric structures of objects while minimizing the impact of visual appearance, thereby reducing redundant details irrelevant to manipulation. Within hybrid action control, imitation learning is combined with visual servo control to manage the moving dynamics of objects. Additionally, an efficient memory modeling strategy is proposed in the imitation learning algorithm to alleviate action prediction ambiguity.
@@
In experiments, we implement a single GEM to support four common skills in dynamic object manipulation, i.e., Pick, Put, Rotate, and Insert. Extensive experiments are conducted in both simulators and the real world to validate the effectiveness of GEM. The experiments suggest that GEM presents promising generalization across diverse environment backgrounds, robot embodiments, motion dynamics, object geometries, and in sim-to-real settings. Especially, as illustrated in Fig. 1, GEM has been successfully deployed in a real canteen to perform tableware collection. Without using demonstrations collected in this canteen, GEM achieves a 97.2% success rate over seven days of testing and more than 10,000 operations.
Summarily, the contributions of this work are as follows:
1) We develop an entropy-based framework to analyze the optimization of imitation learning, guiding the design of an efficient manipulation system with limited demonstrations.
2) We propose GEM, a system combining object-centric geometry perception and hybrid action control to achieve strong generalization in dynamic object manipulation.
3) We demonstrate the first real-world application of imitation learning in dynamic object manipulation, achieving over 97% success rate without in-scene demonstrations.
@@ II-A Imitation Learning for Manipulation
Traditional robotic manipulation methods mostly rely on accurate modeling of the environment to plan action trajectories of robot arms [5]. However, manipulation often involves contact-rich interactions with objects, and accurately modeling these interactions is difficult due to the highly variable nature of contact forces [14]. Moreover, every manipulation task often requires a unique set of models and parameters, making these traditional methods difficult to generalize across different objects, environments, and robot embodiments. This necessitates substantial engineering efforts to adapt the system to new tasks or scenarios [15].
Given these limitations of traditional methods, there is a growing interest in leveraging imitation learning to address the challenges of robotic manipulation [7]. Bypassing the need for explicit system modeling and complex parameter tuning, imitation learning learns manipulation policies directly from expert demonstrations. Specifically, imitation learning trains a policy to predict expert actions based on the observations provided in demonstrations, enabling a robot to mimic expert behavior in similar scenarios. The policy is often implemented as a deep learning network [16], which can capture the complex patterns in demonstrations [17].
The two most commonly employed observation modalities in robotic manipulation are images and 3D points [5]. Correspondingly, existing imitation learning policies can be broadly categorized into image-based [18] and point-based approaches [19]. Images are easy to obtain, and image-based policies can readily leverage large-scale public data for training [20]. However, robotic manipulation mainly concerns the geometric structures of objects, while images primarily capture the appearance of scenes, making it challenging to predict appropriate manipulation actions based solely on visual appearance. In contrast, point clouds directly represent the geometric structures of scenes, and thus they are more suitable for regressing manipulation actions [21]. Unfortunately, point cloud data is relatively scarce compared with images, resulting in a lack of promising point cloud feature extractors [22]. This limitation has historically constrained the performance of point-based manipulation policies [23]. Thankfully, some recent works about large-scale point cloud pre-training have gradually addressed this obstacle [24, 25].
@@ II-B Data Scaling in Manipulation
The rapid progress in fields like large language models [26] suggests that increasing the quantity and diversity of training data is important for enhancing the generalization of deep learning models. Drawing on this insight, the robotic manipulation community has invested significant effort into constructing large-scale demonstration datasets [9, 10]. Leveraging these datasets, numerous studies design algorithms for training end-to-end manipulation policies, with the expectation that these policies exhibit robust generalization across diverse scenarios [11, 27]. Furthermore, some works incorporate additional types of data like object detection datasets to further enrich the knowledge base of manipulation policies [12, 28]. Nevertheless, though much resource has been invested, the generalization capabilities of the obtained policies in unseen environments are still limited.
@@ II-C Sim-to-Real Generalization in Manipulation
Collecting high-quality demonstration data in the real world is labor-intensive. To address this issue, many researchers have explored generating demonstrations in simulated environments [30]. These efforts aim to develop policies trained using only simulated demonstrations yet can generalize to real-world applications. A key first step towards this goal is the development of realistic simulators. Through substantial engineering efforts, several promising simulators have been built, such as MuJuCo [31], Isaac Gym [32], and Isaac Sim [33]. Leveraging these simulators, prior works have used reinforcement learning [34] or predefined kinematic procedures [35] to generate demonstrations. However, reinforcement learning often struggles to discover successful manipulation trajectories due to its inefficient exploration [36]. Grasping objects via predefined kinematic procedures typically requires customized workflows for different scenarios, which is time-consuming. Additionally, current simulators have limited environmental diversity, restricting the variety of generated data.
Another line of works focuses on modeling the environment dynamics of manipulation tasks using classic mechanics and then employs optimization techniques to calculate actions that satisfy task constraints [14]. The observation and action pairs can then be utilized to train deep neural networks [37, 38]. This strategy effectively addresses the slow speed of optimization by distilling the action search process into a more efficient neural network [39]. Nevertheless, for complex manipulation tasks, the computation burden is too heavy to search for satisfactory actions, making it difficult to generate effective demonstrations. To tackle this problem, some works collect demonstrations in simulators through teleoperating simulated robots [40]. Although this method still requires significant human effort, it reduces hardware setup costs compared to real-world data collection.
III Entropy Reduction Theoretical Framework
This section develops a theoretical framework by analyzing the information entropy of imitation learning, providing logical support for developing a generalizable manipulation system with only a few demonstrations.
III-A Robotic Manipulation Problem Formulation
@@
In a specific task target $\tau\in\mathcal{T}$ of robotic manipulation, there usually exist numerous accurate action trajectories that can lead the robot from the initial state $s_{0}$ to the desired goal state $\iota$. We denote all the accurate trajectories as a set $\mathcal{C}$ and each trajectory is represented as a sequence of states and actions $\{s_{0},a_{0},\cdots,s_{T-1},a_{T-1},s_{T}\}$, where $s_{T}=\iota$.
@@ III-B Entropy Analysis of Imitation Learning
Many recent manipulation works derive $\pi$ based on imitation learning [8], which is typically formulated as a supervised learning problem. Imitation learning requires access to a set of expert demonstrations as training data. Assume there are $N_{d}$ demonstration sequences, the expert demonstration set is denoted as $\mathcal{C}_{d}=\{o_{0:T}^{i},\tilde{a}_{0:T}^{i}\}_{i=1}^{N_{d}}$, where $o_{t}^{i}$ and $\tilde{a}_{t}^{i}$ are the observation and expert action at timestamp $t$ of the $i_{\rm th}$ demonstration. The core idea of imitation learning is employing $\mathcal{C}_{d}$ as a surrogate of $\mathcal{C}$. Thus, the objective of imitation learning can be expressed as maximizing the likelihood of expert demonstrations under the policy-induced distribution:

### v2

# Sim-to-Real Dynamic Object Manipulation on Conveyor Systems via Optimization Path Shaping

@@ Abstract
Realizing generalizable dynamic object manipulation on conveyor systems is important for enhancing manufacturing efficiency, as it eliminates specialized engineering for different scenarios. To this end, imitation learning emerges as a promising paradigm, leveraging expert demonstrations to teach a policy manipulation skills. Although the generalization of an imitation learning policy can be improved by increasing demonstrations, demonstration collection is labor-intensive. Besides, public dynamic object manipulation data is scarce. In this work, we address this data scarcity problem via generating demonstrations in a simulator. A significant challenge of using simulated data lies in the appearance gap between simulated and real-world observations. To tackle this challenge, we propose Geometry-Enhanced Model (GEM), which employs our designed appearance noise annealing strategy to shape the policy optimization path, thereby prioritizing the geometry information in observations. Extensive experiments in simulated and real-world tasks demonstrate that GEM can generalize across environment backgrounds, robot embodiments, motion dynamics, and object geometries. Notably, GEM is deployed in a real canteen for tableware collection. Without test-scene data, GEM achieves a success rate of over 97% across more than 10,000 operations.
@@ I Introduction
Robotic manipulation of dynamic objects has broad applications in industrial manufacturing, particularly in conveyor-based assembly lines where robots handle transported products [38, 37]. Compared with the extensively studied static object manipulation problem [23, 18], dynamic object manipulation on conveyor systems poses more challenges due to its demand for more time-efficient motion planning and stricter operational precision. Addressing these challenges is important for enhancing the efficiency and flexibility of industrial automation systems, thereby enabling more economical manufacturing and sustainable production [43].
@@
Therefore, although realizing generalizable dynamic object manipulation on conveyor systems is important, collecting demonstrations is costly, and public data resource is scarce [58]. Given these facts, this work studies whether we can generate demonstrations in a simulator and achieve sim-to-real dynamic object manipulation. Since all real environment states are available in a simulator, we can generate numerous demonstrations via motion planning efficiently. Built upon this idea, we build a dynamic object manipulation simulator based on Isaac Gym [31], as visualized in Fig. 1. Nevertheless, although the simulator handles the collision dynamics between objects well, the appearance gap between simulated and real-world environments is significant.
@@
Figure 1: Primarily using demonstrations collected from a simulator, our method can generalize across diverse environment backgrounds, robot embodiments, motion dynamics, and object geometries. Our method has been reliably deployed in a real canteen to conduct tableware collection. Without using demonstrations collected in this canteen, our method achieves a success rate of over 97% in seven consecutive days of operation, performing more than 10,000 tableware collection operations.
To bridge this appearance gap, our core insight is that the information in policy observation can be decomposed into geometric structure and visual appearance. Importantly, manipulation primarily relies on geometric structure, which is consistent between simulated and real-world environments. Based on this insight, we propose a policy named Geometry-Enhanced Model (GEM), which takes colored 3D points from RGB-D cameras as input. In GEM, we propose an appearance noise annealing strategy to shape the policy optimization path, reducing the reliance of the policy on appearance while enhancing its focus on geometry. In this way, sim-to-real generalization is enabled. Moreover, we propose a control strategy that disentangles manipulation into tracking actions and interaction actions. This strategy empowers GEM to manipulate objects moving in unseen velocities.
@@
We conduct extensive experiments in both simulated and real-world tasks to validate the effectiveness of GEM. The experiments suggest that GEM presents promising generalization across diverse environment backgrounds, robot embodiments, motion dynamics, and object geometries. Especially, as illustrated in Fig. 1, GEM has been successfully deployed in a real canteen to collect tableware. Without using demonstrations collected in this canteen, GEM achieves a 97.2% success rate over seven days of testing and more than 10,000 operations.
Summarily, the technical contributions of this work are summarized as follows:
1) We build the first simulation environment for generating dynamic object manipulation data on conveyor systems.
2) We propose a neural policy GEM, which realizes sim-to-real generalization by prioritizing the exploitation of geometry information while reducing the reliance on appearance.
3) We design a control strategy that disentangles manipulation into tracking and interaction actions, enabling GEM to manipulate dynamic objects of diverse motion patterns.
@@ II-A Imitation Learning for Manipulation
Traditional robotic manipulation methods mostly rely on accurate modeling of the environment to plan action trajectories of robot arms [46]. However, manipulation often involves contact-rich interactions with objects, and accurately modeling these interactions is difficult due to the highly variable nature of contact forces [19]. Moreover, every single manipulation task often requires a unique set of models and parameters, making these traditional methods difficult to generalize across different objects, environments, and robot embodiments. This necessitates substantial engineering efforts to adapt the system to new tasks or scenarios [59].
Given these limitations of traditional methods, there is a growing interest in leveraging imitation learning to address the challenges of robotic manipulation [55]. Bypassing the need for explicit system modeling and complex parameter tuning, imitation learning learns manipulation policies directly from expert demonstrations. Specifically, imitation learning trains a policy to predict expert actions based on the observations provided in demonstrations, enabling a robot to mimic expert behavior in similar scenarios. The policy is often implemented as a deep learning network [26], which can capture the complex patterns of demonstrations [6].
The two most commonly employed observation modalities in robotic manipulation are images and 3D points [46]. Correspondingly, existing imitation learning policies can be broadly categorized into image-based [13] and point-based approaches [44]. Images are easy to obtain, and image-based policies can readily leverage large-scale public data for training [33]. However, robotic manipulation mainly concerns the geometric structures of objects, while images primarily capture the appearance of scenes, making it challenging to predict appropriate manipulation actions solely based on visual appearance. In contrast, point clouds directly represent the geometric structures of scenes, and thus they are more suitable for estimating manipulation actions [52]. Unfortunately, point cloud data is relatively scarce compared with images, resulting in a lack of promising point cloud feature extractors [39]. This limitation has historically constrained the performance of point-based manipulation policies [15]. Thankfully, some recent works about large-scale point cloud pre-training have gradually addressed this obstacle [53, 51].
@@ II-B Data Scaling in Manipulation
The rapid progress in fields like large language models [1] suggests that increasing the quantity and diversity of training data is important for enhancing the generalization of deep learning models. Drawing on this insight, the robotic manipulation community has invested significant effort into constructing large-scale demonstration datasets [35, 22]. Leveraging these datasets, numerous studies design algorithms for training end-to-end manipulation policies, with the expectation that these policies exhibit robust generalization across diverse scenarios [24, 29]. Furthermore, some works incorporate additional types of data like object detection datasets to further enrich the knowledge of manipulation policies [4, 3]. Nevertheless, though much resource has been invested, the generalization capabilities of the obtained policies in unseen environments are still limited.
@@ II-C Sim-to-Real Generalization in Manipulation
Collecting high-quality demonstration data in the real world is labor-intensive. To address this obstacle, many researchers have explored generating demonstrations in simulated environments [32]. These efforts aim to develop policies that solely employ simulated demonstrations to train yet can generalize to real-world applications. A key step towards this goal is the development of realistic simulators. Through substantial engineering efforts, several promising simulators have been built, such as MuJuCo [48], Isaac Gym [31], and Isaac Sim [34]. Leveraging these simulators, prior works have used reinforcement learning [20] or predefined kinematic procedures [9] to generate demonstrations. However, reinforcement learning often struggles to discover appropriate manipulation trajectories due to its inefficient exploration [47]. Grasping objects via kinematic procedures typically requires customized workflows for different scenarios, which is time-consuming. Additionally, current simulators have limited environmental diversity, restricting the variety of the generated data.
Another line of literature focuses on modeling the environment dynamics of manipulation tasks using classic mechanics and then employs optimization techniques to calculate actions that satisfy task constraints [19]. The observation and action pairs can then be utilized to train deep neural networks [12, 49]. This strategy effectively addresses the slow speed of optimization by distilling the action search process into a more efficient neural network [16]. Nevertheless, for complex manipulation tasks, the computation burden for searching promising action trajectories could be heavy, making it difficult to generate effective demonstrations. To tackle this problem, some works collect demonstrations in simulators through teleoperating robots [61]. Although this method still requires significant human effort, it reduces hardware setup costs compared to real-world data collection.
III Problem Analysis
This section outlines the policy design principle of this work, providing a theoretical analysis of how we shape the optimization path of an imitation learning policy. Specifically, Section III-A formulates the policy optimization objective of robotic manipulation. Next, Section III-B discusses policy observation information and explains why 3D geometry is more important than 2D appearance in embodied tasks. Subsequently, Section III-C introduces the core process of imitation learning. Finally, Section III-D elaborates the concept of optimization path shaping in imitation learning, which is the theoretical foundation of this work.
III-A Problem Formulation
@@
In a specific task target $\tau\in\mathcal{T}$ of robotic manipulation, there usually exist numerous accurate action trajectories that can lead the robot from the initial state $s_{0}$ to the desired goal state $\iota$. We denote all the accurate trajectories as a set $\mathcal{C}$ and each trajectory can be represented as a sequence of states and actions $\{s_{0},a_{0},\cdots,s_{T-1},a_{T-1},s_{T}\}$, where $s_{T}=\iota$.
@@
Figure 2: The 3D geometry of objects usually plays a more important role than 2D appearance in scheduling manipulation actions. As shown, although the left and middle objects share similar color and texture, the appropriate gripper poses to pick them up are different. By contrast, though the middle and right objects belong to different object categories with distinct appearances, they can be picked up with a similar gripper pose.
III-B Policy Observation Analysis
According to the aforementioned formulation, the environment observation $o_{t}$ provides the information that is utilized by $\pi$ to schedule actions. The two most commonly employed observation modalities for recent manipulation policies are images captured by cameras [13] and point clouds returned by depth sensors [56]. Images directly describe environment appearance, allowing $\pi$ to more accurately identify object categories and materials. Point clouds focus more on object geometry, enabling a policy to determine manipulation poses more precisely. Both the appearance and geometry information are valuable for successful manipulation. To combine their benefits, this work adopts colored point clouds obtained from RGB-D cameras as policy input, enabling the policy to perceive both appearance and geometry information appropriately.
Notably, in the tasks of dynamic object manipulation on conveyor belts, geometry is more critical than appearance. Without color information, a policy can still identify the moving objects on a conveyor belt in most cases. Moreover, the policy can often determine accurate manipulation actions based only on geometry. For instance, as illustrated in Fig. 2, while the left and middle objects share similar color and texture, their accurate gripper poses for grasping are different. In contrast, despite belonging to distinct object categories with unique appearances, the middle and right objects can be picked up using similar gripper poses. Therefore, a policy should focus more on exploiting the geometry information.
In addition, as visualized in Fig. 1, while existing simulators can effectively replicate the geometric structure of objects, the rendered appearance still exhibits a significant gap compared to the real world. Thankfully, as indicated by the above analysis, successful manipulation in our explored tasks primarily requires accurate geometry information. Hence, in this work, we expect $\pi$ to concentrate on 3D geometry rather than 2D appearance. To achieve this, we design strategies to shape the optimization dynamics in imitation learning. For a clear explanation of these strategies, we first introduce the core process of imitation learning in Section III-C.
III-C Imitation Learning
Imitation learning is typically formulated as a supervised learning problem [60]. It requires access to a set of expert demonstrations as training data. Assume there are $N_{d}$ demonstration sequences, the expert demonstration set is denoted as $\mathcal{C}_{d}=\{o_{0:T}^{i},\tilde{a}_{0:T}^{i}\}_{i=1}^{N_{d}}$, where $o_{t}^{i}$ and $\tilde{a}_{t}^{i}$ are the observation and expert action at timestamp $t$ of the $i_{\rm th}$ demonstration. The core idea of imitation learning is employing $\mathcal{C}_{d}$ as a surrogate of the successful manipulation set $\mathcal{C}$. Thus, the objective of imitation learning can be expressed as maximizing the likelihood of expert demonstrations under the policy-induced distribution: